# Hurwitz moduli varieties parameterizing pointed covers of an algebraic curve with a fixed monodromy group

### v3

@@ 4. Universal families of pointed covers with fixed monodromy group
The graph $\Gamma$ of $\psi$ is contained in the set-theoretical fiber product $X\times_{Y\times S}\mathcal{X}(y_{0},\lambda_{0})_{S}$, which is a Zariski closed subset of $X\times\mathcal{X}(y_{0},\lambda_{0})_{S}$, so it contains the closure $\overline{\Gamma}$. Therefore the projection morphism $\overline{\Gamma}\to X$ has finite fibers. Applying [20, Theorem 2] one concludes that $\psi$ can be extended to an $S$-morphism $\tilde{\psi}:X\to\mathcal{X}(y_{0},\lambda_{0})_{S}$. For every $s\in S$ the composition $X_{s}\overset{\tilde{\psi}_{s}}{\longrightarrow}(\mathcal{X}(y_{0},\lambda_{0})_{S})_{s}\overset{\sim}{\longrightarrow}\mathcal{X}(y_{0},\lambda_{0})_{u(s)}$ is a morphism whose restriction on $X^{\prime}_{s}$ coincides with $h^{\prime}_{s}$. Hence this composition equals $h_{s}$. This implies that $h$ equals the composition of morphisms $X\overset{\tilde{\psi}}{\longrightarrow}\mathcal{X}(y_{0},\lambda_{0})_{S}\longrightarrow\mathcal{X}(y_{0},\lambda_{0})$, so $h:X\to\mathcal{X}(y_{0},\lambda_{0})$ is a morphism. Furthermore $\tilde{\psi}_{s}$ is an isomorphism for every $s\in S$. Applying Lemma 4.16 to the smooth, proper morphisms $X\to S$ and $\mathcal{X}(y_{0},\lambda_{0})_{S}\to S$ one concludes that $\tilde{\psi}:X\to\mathcal{X}(y_{0},\lambda_{0})_{S}$ is an isomorphism, so Diagram (27) is Cartesian. One has that $\phi_{S}\circ\tilde{\psi}=f$ and $\tilde{\psi}(\eta(s))=\psi(\eta(s))=\xi_{S}(s)$ for $\forall s\in S$, since $h(\eta(s))=\xi(u(s))$ by the construction of the map $h$. Therefore $(f:X\to Y\times S,\eta)$ is equivalent to the pullback of the family (26) by the morphism $u:S\to H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$. ∎
@@
(29) ${T\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{u}}$$\scriptstyle{\mu}$${H^{G}_{n}(Y,y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u}$${H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})}$
The morphism $\mu:T\to S$ is an étale Galois cover with Galois group $N(\lambda_{0})$ (cf.[22, Prop. A.7.1.3]). Let $(f_{T}:X_{T}\to Y\times T,\eta_{T})$ be the pullback of $(f:X\to Y\times S,\eta)$ by $\mu:T\to S$ (cf. § 3.8). One has the following commutative diagram of morphisms in which the two squares are Cartesian
@@
Therefore the composed diagram is Cartesian as well [11, Proposition 4.16]. Let $(\varphi:\mathcal{X}(y_{0})\to Y\times H^{G}_{n}(Y,y_{0}),\theta:H^{G}_{n}(Y,y_{0})\to\mathcal{X}(y_{0}))$ be the family of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ defined in Proposition 4.6. Let $(f_{1}:X_{1}\to Y\times T,\eta_{1})$ be its pullback by $\tilde{u}:T\to H^{G}_{n}(Y,y_{0})$. One has the following commutative diagram of morphisms in which the left square is Cartesian and the right one is (23)
@@
Let $(p_{T}:\mathcal{C}\to Y\times T,\zeta_{T})$ be the pullback of $(p:\mathcal{C}(y_{0})\to Y\times H^{G}_{n}(Y,y_{0}),\zeta)$ by $\tilde{u}:T\to H^{G}_{n}(Y,y_{0})$ [21, § 5.2]. One has that $\mathcal{C}$ is reduced and may be identified with the closed subvariety $\mathcal{C}=\{(z,t)|\pi_{2}\circ p(z)=\tilde{u}(t)\}$ of $\mathcal{C}(y_{0})\times T$. The induced action of $G$ is defined by $g(z,t)=(gz,t)$ and $\zeta_{T}(t)=(\zeta(\tilde{u}(t)),t)$ for $\forall t\in T$. Let $j:\mathcal{C}\to\mathcal{C}(y_{0})$ be the $G$-equivariant morphism $j(z,t)=z$. Let $H=G(\lambda_{0})$ and let $\rho:\mathcal{C}(y_{0})\to\mathcal{X}(y_{0})$ be the $H$-invariant morphism defined by $\rho(z)=\pi(\lambda_{0},z)$. Let $\rho_{1}:\mathcal{C}\to X_{1}$ be the morphism defined by $\rho_{1}(z,t)=(\rho(z),t)$. It is $H$-invariant and fits in the following commutative diagram of morphisms.
@@
One has that $f_{1}\circ\rho_{1}=p_{T}$ and $\varphi\circ\rho=p$, so the composed diagram is Cartesian. The right square is Cartesian, so the left square is Cartesian as well [11, Proposition 4.16]. By Proposition 4.7 $\rho$ induces an isomorphism $\overline{\rho}:\mathcal{C}(y_{0})/H\overset{\sim}{\longrightarrow}\mathcal{X}(y_{0})$. The formation of quotients by $H$ commutes with base change [22, Prop. A.7.1.3], so $\rho_{1}$ induces an isomorphism $\overline{\rho}_{1}:\mathcal{C}/H\to X_{1}$ over $Y\times T$. Furthermore $\rho_{1}$ transforms the section $\zeta_{T}:T\to\mathcal{C}$ into $\eta_{1}:T\to X_{1}$. In fact, for every $t\in T$ one has
@@
(33) ${T\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{u}}$$\scriptstyle{\mu}$${H^{G}_{\underline{n}}(Y,y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u}$${H^{\Lambda,G}_{\underline{n},\lambda_{0}}(Y,y_{0})}$
and using the universal family of pointed $G$-covers of $(Y,y_{0})$ of branching type $\underline{n}$, $(p_{\underline{n}}:\mathcal{C}_{\underline{n}}(y_{0})\to Y\times H^{G}_{\underline{n}}(Y,y_{0}),\zeta_{\underline{n}}$) (cf. [21, Theorem 5.8]). ∎
@@
be the smooth, proper family of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ branched in $n$ points associated with $\lambda_{1}\in\Lambda$ (cf. Theorem 4.13). Then there exists an isomorphism $h:\mathcal{X}(y_{0},\lambda_{1})\to\mathcal{X}(y_{0},\lambda_{0})$ which fits in the following diagram
(35) ${\mathcal{X}(y_{0},\lambda_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$$\scriptstyle{\phi_{1}}$${\mathcal{X}(y_{0},\lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${Y\times H^{\Lambda,G}_{n,\lambda_{1}}(Y,y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{id\times u}$${Y\times H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})}$
@@
Acknowledgments.
The author is associate member of the Institute of Mathematics and Informatics of the Bulgarian Academy of Sciences.

### v4

@@ 4. Universal families of pointed covers with fixed monodromy group
The graph $\Gamma$ of $\psi$ is contained in the set-theoretical fiber product $X\times_{Y\times S}\mathcal{X}(y_{0},\lambda_{0})_{S}$, which is a Zariski closed subset of $X\times\mathcal{X}(y_{0},\lambda_{0})_{S}$, so it contains the closure $\overline{\Gamma}$. Therefore the projection morphism $\overline{\Gamma}\to X$ has finite fibers. Applying [21, Theorem 2] one concludes that $\psi$ can be extended to an $S$-morphism $\tilde{\psi}:X\to\mathcal{X}(y_{0},\lambda_{0})_{S}$. For every $s\in S$ the composition $X_{s}\overset{\tilde{\psi}_{s}}{\longrightarrow}(\mathcal{X}(y_{0},\lambda_{0})_{S})_{s}\overset{\sim}{\longrightarrow}\mathcal{X}(y_{0},\lambda_{0})_{u(s)}$ is a morphism whose restriction on $X^{\prime}_{s}$ coincides with $h^{\prime}_{s}$. Hence this composition equals $h_{s}$. This implies that $h$ equals the composition of morphisms $X\overset{\tilde{\psi}}{\longrightarrow}\mathcal{X}(y_{0},\lambda_{0})_{S}\longrightarrow\mathcal{X}(y_{0},\lambda_{0})$, so $h:X\to\mathcal{X}(y_{0},\lambda_{0})$ is a morphism. Furthermore $\tilde{\psi}_{s}$ is an isomorphism for every $s\in S$. Applying Lemma 4.16 to the smooth, proper morphisms $X\to S$ and $\mathcal{X}(y_{0},\lambda_{0})_{S}\to S$ one concludes that $\tilde{\psi}:X\to\mathcal{X}(y_{0},\lambda_{0})_{S}$ is an isomorphism, so Diagram (27) is Cartesian. One has that $\phi_{S}\circ\tilde{\psi}=f$ and $\tilde{\psi}(\eta(s))=\psi(\eta(s))=\xi_{S}(s)$ for $\forall s\in S$, since $h(\eta(s))=\xi(u(s))$ by the construction of the map $h$. Therefore $(f:X\to Y\times S,\eta)$ is equivalent to the pullback of the family (26) by the morphism $u:S\to H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$. ∎
@@
The morphism $\mu:T\to S$ is an étale Galois cover with Galois group $N(\lambda_{0})$ (cf.[23, Prop. A.7.1.3]). Let $(f_{T}:X_{T}\to Y\times T,\eta_{T})$ be the pullback of $(f:X\to Y\times S,\eta)$ by $\mu:T\to S$ (cf. §3.8). One has the following commutative diagram of morphisms in which the two squares are Cartesian:
@@
Therefore the composed diagram is Cartesian as well [11, Proposition 4.16]. Let $(\varphi:\mathcal{X}(y_{0})\to Y\times H^{G}_{n}(Y,y_{0}),\theta:H^{G}_{n}(Y,y_{0})\to\mathcal{X}(y_{0}))$ be the family of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ defined in Proposition 4.6. Let $(f_{1}:X_{1}\to Y\times T,\eta_{1})$ be its pullback by $\tilde{u}:T\to H^{G}_{n}(Y,y_{0})$. One has the following commutative diagram of morphisms in which the left square is Cartesian and the right one is (23):
(31) ${X_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h_{1}}$$\scriptstyle{f_{1}}$${\mathcal{X}(y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$$\scriptstyle{\kappa}$${\mathcal{X}(y_{0},\lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${Y\times T\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{id\times\tilde{u}}$${Y\times H^{G}_{n}(Y,y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{id\times\nu}$${Y\times H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0}).}$
@@
Let $(p_{T}:\mathcal{C}\to Y\times T,\zeta_{T})$ be the pullback of $(p:\mathcal{C}(y_{0})\to Y\times H^{G}_{n}(Y,y_{0}),\zeta)$ by $\tilde{u}:T\to H^{G}_{n}(Y,y_{0})$ [22, §5.2]. One has that $\mathcal{C}$ is reduced and may be identified with the closed subvariety $\mathcal{C}=\{(z,t)|\pi_{2}\circ p(z)=\tilde{u}(t)\}$ of $\mathcal{C}(y_{0})\times T$. The induced action of $G$ is defined by $g(z,t)=(gz,t)$ and $\zeta_{T}(t)=(\zeta(\tilde{u}(t)),t)$ for $\forall t\in T$. Let $j:\mathcal{C}\to\mathcal{C}(y_{0})$ be the $G$-equivariant morphism $j(z,t)=z$. Let $H=G(\lambda_{0})$ and let $\rho:\mathcal{C}(y_{0})\to\mathcal{X}(y_{0})$ be the $H$-invariant morphism defined by $\rho(z)=\pi(\lambda_{0},z)$. Let $\rho_{1}:\mathcal{C}\to X_{1}$ be the morphism defined by $\rho_{1}(z,t)=(\rho(z),t)$. It is $H$-invariant and fits in the following commutative diagram of morphisms:
@@
One has that $f_{1}\circ\rho_{1}=p_{T}$ and $\varphi\circ\rho=p$, so the composed diagram is Cartesian. The right square is Cartesian, so the left square is Cartesian as well [11, Proposition 4.16]. By Proposition 4.7 $\rho$ induces an isomorphism $\overline{\rho}:\mathcal{C}(y_{0})/H\overset{\sim}{\longrightarrow}\mathcal{X}(y_{0})$. The formation of quotients by $H$ commutes with base change [23, Prop. A.7.1.3], so $\rho_{1}$ induces an isomorphism $\overline{\rho}_{1}:\mathcal{C}/H\to X_{1}$ over $Y\times T$. Furthermore $\rho_{1}$ transforms the section $\zeta_{T}:T\to\mathcal{C}$ into $\eta_{1}:T\to X_{1}$. In fact, for every $t\in T$ one has
@@
[…]
and using the universal family of pointed $G$-covers of $(Y,y_{0})$ of branching type $\underline{n}$, $(p_{\underline{n}}:\mathcal{C}_{\underline{n}}(y_{0})\to Y\times H^{G}_{\underline{n}}(Y,y_{0}),\zeta_{\underline{n}}$) (cf. [22, Theorem 5.8]). ∎
@@
be the smooth, proper family of pointed $(\Lambda,G)$-covers of $(Y,y_{0})$ branched in $n$ points associated with $\lambda_{1}\in\Lambda$ (cf. Theorem 4.13). Then there exists an isomorphism $h:\mathcal{X}(y_{0},\lambda_{1})\to\mathcal{X}(y_{0},\lambda_{0})$ which fits in the following diagram:
(35) ${\mathcal{X}(y_{0},\lambda_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$$\scriptstyle{\phi_{1}}$${\mathcal{X}(y_{0},\lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${Y\times H^{\Lambda,G}_{n,\lambda_{1}}(Y,y_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{id\times u}$${Y\times H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0}),}$